# Clock Auctions Augmented with Unreliable Advice

### v1

@@ A.1 Proof of Lemma 2
$\sum_{k^{\prime}=1}^{n}{\frac{R_{t}}{k^{\prime}}}=R_{t}\cdot H_{n}.\qed$
@@ A.2 Proof of Lemma 6
$\sum_{t^{\prime}=1}^{t}{R_{i^{\prime}}\cdot H_{n}}=H_{n}\sum_{t^{\prime}=1}^{t}{10^{t^{\prime}}}=H_{n}\cdot\frac{10^{t+1}-1}{9}\leq R_{t}\cdot\frac{10H_{n}}{9}\qed$
@@ D.2 Proof of Corollary 1
$\sum_{t=1}^{t^{\prime}}{\frac{\beta}{2}R^{P}_{t-1}}=\sum_{t=1}^{t^{\prime}-1}{\frac{\beta}{2}R_{t-1}^{P}}+\frac{\beta}{2}R^{P}_{t^{\prime}-1}\leq\beta R_{t^{\prime}-1}^{P}.\qed$
@@ D.3 Proof of Lemma 11
Observe that when we terminate, by assumption that we output a subset of the predicted set, we did so because we did not satisfy either condition in Line 4. But in the previous iteration of the while loop we have that the revenue of the predicted set was $(\alpha-1)$ times the total welfare rejected from the predicted set. But then the social welfare remaining in the predicted set is at least a $\alpha$-factor of the total social welfare in the predicted set. ∎

### v2

@@ A.1 Proof of Lemma 2
$\sum_{k^{\prime}=1}^{n}{\frac{R_{t}}{k^{\prime}}}=R_{t}\cdot H_{n},$
completing the proof. ∎
@@ A.2 Proof of Lemma 6
$\sum_{t^{\prime}=1}^{t}{R_{i^{\prime}}\cdot H_{n}}=H_{n}\sum_{t^{\prime}=1}^{t}{10^{t^{\prime}}}=H_{n}\cdot\frac{10^{t+1}-1}{9}\leq R_{t}\cdot\frac{10H_{n}}{9},$
as desired. ∎
A.3 Proof of Corollary 1
Proof.
Suppose we have an algorithm which gives a $\rho$-approximation to the underlying algorithmic problem of maximizing welfare/revenue on a fixed set of values. We first argue that the consistency of FollowTheUnpredictedLeader when using this approximation algorithm (in place of an exact algorithm) remains $(1+\epsilon)$. To see this, observe that we can exactly compute the revenue within the predicted set in every round in polynomial time since we can simply sum the prices of the bidders in the predicted set. As such, the arguments in Lemma 7 still hold when we use an approximation algorithm in Lines 7 and 9 of FollowTheUnpredictedLeader on the unpredicted sets.
It remains to verify that the robustness of FollowTheUnpredictedLeader is $O(\rho\cdot\frac{1}{\epsilon}\log{n})$. First observe that, by definition of the fact that we are using a $\rho$-approximate algorithm in Line 7, the maximum welfare rejected from any unpredicted set in iteration $t$ of the while loop of Mechanism 2 at the end of line 7 is $\rho\cdot 2R_{t}\gamma H_{n}$. Similarly, by the same reasoning as in the proof of Lemma 2, the total welfare rejected from any unpredicted set in line 9 of Mechanism 2 in iteration $t$ is at most $\rho R_{t}\cdot H_{n}$ (since $k$ bidders in any set accepting a price $\rho R_{t}/k$ would mean that the best revenue is weakly higher than $\rho R_{t}$ and a $\rho$-approximate algorithm would find a set with revenue at least $R_{t}$). Combining these two gives that the total welfare rejected from any unpredicted set in any iteration $t$ is at most $R_{t}\cdot(2\gamma+1)\rho H_{n}$. As in Lemma 3, summing over all rounds then gives that the total welfare rejected from any unpredicted set from rounds $1$ to $t$ is at most $R_{t}\cdot\frac{(2\gamma+1)10\rho H_{n}}{9}$.
Following similar reasoning to Lemmas 3 and 5, we consider cases depending on whether a predicted or unpredicted set wins. Let $\hat{t}$ denote the final iteration of the while loop of Mechanism 2. If Mechanism 2 outputs a subset of the predicted set, observe that the revenue of the predicted bidders is at least $R_{\hat{t}-1}=\frac{R_{\hat{t}}}{10}$. But then, since the total welfare in any unpredicted set is at most $R_{\hat{t}}\cdot\frac{(2\gamma+1)10\rho H_{n}}{9}$ (since we rejected all the bidders in some round between $1$ and $\hat{t}$), we obtain that Mechanism 2 outputs a set with at least a $\frac{9}{\rho\cdot(2\gamma+1)100H_{n}}$-fraction of the optimal welfare. Suppose, instead, that Mechanism 2 outputs a subset of some unpredicted set (i.e., we reach line 2) in the final iteration of the while loop $\hat{t}$. By the same arguments in Lemma 4, we know that the revenue among active bidders in the highest revenue set is $R_{\hat{t}}$ when we reach line 2. As above, the total welfare rejected from any unpredicted set by this point is at most $R_{\hat{t}}\cdot\frac{(2\gamma+1)10\rho H_{n}}{9}$. But then, we have that the revenue among active bidders in the highest revenue set when we reach line 2 is a $\frac{9}{\rho\cdot 10(2\gamma+1)H_{n}}$-fraction of the maximum social welfare from any feasible set of rejected unpredicted bidders. We can then apply similar reasoning to the proof of Lemma 5 to complete the proof. As noted in [CGS22], the WFCA can also readily admit the use of a $\rho$-approximation for maximizing the revenue at a fixed set of clock prices. In this case, the approximation guarantee given the WFCA is $\rho\cdot 2H_{n}$ [CGS22, Theorem 5]. Moreover, the WFCA remains revenue monotone even when using an approximation algorithm. Combining these observations with our above reasoning, we have that the revenue from the set output by Mechanism 2 is a $\frac{9}{\rho\cdot 10(2\gamma+1)H_{n}}$-fraction of the welfare rejected before running the WFCA and the welfare from the set output by Mechanism 2 is a $\frac{1}{\rho\cdot 2H_{n}}$. Putting this all together, we have that Mechanism 2 with any $\gamma>10/9$ and a $\rho$-approximate black-box achieves at least a $\frac{9}{\rho\cdot 10(2\gamma+1)H_{n}}+\frac{1}{\rho\cdot 2H_{n}}$-fraction of the optimal welfare when the prediction is incorrect and Mechanism 2 outputs a subset of some unpredicted set. Combining our two cases gives that Mechanism 2 is $O(\rho\cdot\frac{1}{\epsilon}\log{n})$-robust, as desired. ∎
@@ D.2 Proof of Corollary 2
$\sum_{t=1}^{t^{\prime}}{\frac{\beta}{2}R^{P}_{t-1}}=\sum_{t=1}^{t^{\prime}-1}{\frac{\beta}{2}R_{t-1}^{P}}+\frac{\beta}{2}R^{P}_{t^{\prime}-1}\leq\beta R_{t^{\prime}-1}^{P},$
as desired. ∎
@@ D.3 Proof of Lemma 11
Observe that when we terminate, by assumption that we output a subset of the predicted set, we did so because we did not satisfy either condition in Line 4. But in the previous iteration of the while loop we have that the revenue of the predicted set was $(\alpha-1)$ times the total welfare rejected from the predicted set. The social welfare remaining in the predicted set is then at least a $\alpha$-factor of the total welfare of the predicted set. ∎